It takes you 8 hours to paint one room by yourself. If you work for x hours, what portion of the job do you complete?
Answer by stanbon(75887)

It takes you 8 hours to paint one room by yourself. If you work for x hours, what portion of the job do you complete?
-----------------
rate = 1/8 job/hr
----------------------
work = rate*time
---
work = (1/8)*x = x/8 job
==============================
Cheers,
Stan H.
===============